# Context Representation via Action-Free Transformer encoder-decoder for Meta Reinforcement Learning

### v1

@@ 5 Experiments
This section presents experimental results for CRAFT belief inference model, built upon a transformer encoder-decoder architecture and designed for adaptive meta-RL dealing with multi-task robotic environments. The experiments evaluate the ability of the model to infer task structure and adapt policies across diverse tasks in a conventional simulation benchmark designed for such an objective. The section is organized into environment setup, implementation details, performance results, qualitative insights, and discussion.
@@ 5.1 Environments and Training Pipeline
Following some isolated update iterations of the RL agent, i.e., actor and critic networks, evaluation is then performed on all variants of both the training and test sets $\mathcal{T}^{\text{train}}\cup\mathcal{T}^{\text{test}}$. Figure˜4 illustrates how the environments are sampled in meta-training and adaptation phases. It should be noted that the RL updates need a separate buffer for keeping all the action and belief information in the interactions. To quantify adaptation to each task, the average return and success rate of all its variations are computed. This separation enables detailed analysis of generalization to familiar and unfamiliar environments under both identity and parametric diversity.
@@ 5.2 Training Implementation Details
    $\mathbf{u}_{t}^{s}=\text{FFN(Multi-Head(CausalSelfAttn))}(f^{s}_{q}(\mathbf{E^{\prime}}^{s}),f^{s}_{k}(\mathbf{E^{\prime}}^{s}),f^{s}_{v}(\mathbf{E^{\prime}}^{s}))$ ;
@@
    $\mathbf{u}_{t}^{r}=\text{FFN(Multi-Head(CausalSelfAttn))}(f^{r}_{q}(\mathbf{E^{\prime}}^{r}),f^{r}_{k}(\mathbf{E^{\prime}}^{r}),f^{r}_{v}(\mathbf{E^{\prime}}^{r}))$ ;
@@
    $h_{t}=\text{FFN(Multi-Head(CausalCrossAttn))}(f^{r}_{q}(\mathbf{u}^{s}),f^{r}_{k}(\mathbf{u}^{s}),f^{r}_{v}(\mathbf{u}^{r}))$ ;
@@ 5.3.1 Quantitative Findings
The quantitative evaluation focuses on two performance metrics, return and success rate, across all task variations in the MetaWorld ML-10 benchmark. The average values of these metrics are plotted against the number of training interaction frames in Figure˜5 and Figure˜6. Results are reported separately for training and test environments, ensuring that generalization can be meaningfully assessed. Table˜2 and Table˜3 present more detailed numerical findings, showing the post-training generalization performance in each environment for CRAFT and other Bayes-adaptive baselines.
@@
During many stages of training, CRAFT achieves higher final-episode return and success compared to the mean over the $n_{H}$ episodes in a meta-episode, denoted as $R_{f}-\overline{R}_{H^{+}}$ and $S_{f}-\overline{S}_{H^{+}}$, respectively. The corresponding values are plotted against the number of meta-training interaction frames in Figures˜7 and 8. As can be seen, this relative advantage remains in the later stages of training for the action-free methods, suggesting that the policy can still benefit from continued exploration and training. This indicates that accumulated experience within the meta-episode contributes meaningfully to improved performance, reflecting effective belief updates in the BAMDP framework. In contrast, other baseline methods show weaker improvement—or even degradation in some cases—between early and final episodes. The ability of transformer-based models to better extract information from long contexts compared to memory-based sequence models contributes to this result. It should also be noted that CRAFT, particularly when a state decoder is included during training, shows the most noticeable benefit during intermediate training stages, when the belief model is still developing and the policy is actively exploring. This further supports the value of gradient isolation in transformer-based context encoder training, which enables informative task inference independently and provides a strong prior for the adaptation process.
@@ 5.3.2 Qualitative Findings
To further analyze the structure of the learned latent space, the pairwise projections of the five-dimensional task belief vectors are plotted. A 2-D plot is generated for each pair of the latent space dimensions, resulting in ten distinct subplots covering all one-vs-one combinations shown in Figure˜9. Some task environments cluster closely in certain latent dimensions while remaining dispersed in others, suggesting that different latent axes are responsible for encoding distinct aspects of task variation. A 3-D projection of the latent vectors is illustrated in Figure˜10. By finding the means and variances of the 3-D UMAP projections computed for the latent representations of each of the 15 tasks [mcinnes2018umap], the estimated distributions are visualized as ellipsoids, indicating the variability region of each task in the latent space.
@@
Figure˜11 shows the average latent vector in all evaluation episodes of all variations of each environment. Some tasks consistently exhibit large values along specific latent dimensions, effectively dominating those axes. For instance, the lever-pull task is represented by a very low value along axis one and a large value on axis three, while the door-open task shows the highest value on axis one compared to all others. In general, conceptually similar tasks tend to share these dominant dimensions, indicating that the model has learned a structured and interpretable belief space. This organization reflects the transformer encoder-decoder’s ability to disentangle task-specific features. CRAFT was able to extract meaningful representations without access to action information, supporting its role as a reliable inference module in adaptive meta-RL.
Figure 11: The numerical values of the average task belief vector across all tasks in the MetaWorld ML-10 benchmark
@@ 7 Future Directions
The former possibility is quite straightforward in terms of implementation. However, the latter needs further explanation. As an example, consider a robotic environment expressed by two different MDPs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ per Equation˜4. These two MDPs share the same action space ($\mathcal{A}$), dynamic functions ($\mathcal{P}_{S},\mathcal{P}_{R}$), discount factor ($\gamma$), and episode length ($H$). However, the two MDPs differ in the state space domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ and optimal policies $\pi^{*}_{1}$ and $\pi^{*}_{2}$, respectively. For instance, $\mathcal{M}_{1}$ can receive state observations in the numerical space from sensor data, versus $\mathcal{M}_{2}$ receiving workspace images as observations. Assuming the essential Markov property holds for both MDPs, it is possible to infer a shared latent domain where $\Psi(\mathcal{S}_{1})\cong\Phi(\mathcal{S}_{2})$, as conceptualized in Figure˜12, such that $\Psi(\mathcal{M}_{1})\simeq\Phi(\mathcal{M}_{2})$. These transformed MDPs can be governed with a shared optimal policy $\pi^{*}\simeq\pi^{*}(\Psi(\cdot))\simeq\pi^{*}(\Phi(\cdot))$.
If learning the mappings $\Psi(\cdot)$ and $\Phi(\cdot)$, and their inverse functions $\Psi^{-1}(\cdot)$ and $\Phi^{-1}(\cdot)$, is tractable, it is possible to train an action-agnostic, and hence agent-agnostic, foundational transformer encoder-decoder from offline data in the source domain and fine-tune it with additional adapters to enable adaptation in the target domain. This means the proposed belief model can be trained on an MDP with numerical state observations. By learning observation adapters $q^{\text{belief}}$ and $q^{\text{policy}}$, and fine-tuning the RL agent consisting of value function $V_{\nu}$ and $\pi_{\zeta}$, the context-adaptive meta-RL pipeline can acclimate to visual state observations online as outlined in Figure˜13.

### v2

@@ 5 Experiments
This section presents experimental results for the CRAFT belief inference model, built upon a transformer encoder-decoder architecture and designed for adaptive meta-RL dealing with multi-task robotic environments. The experiments evaluate the ability of the model to infer task structure and adapt policies across diverse tasks in a conventional simulation benchmark designed for such an objective. The section is organized into environment setup, implementation details, performance results, qualitative insights, and discussion.
@@ 5.1 Environments and Training Pipeline
Following some isolated update iterations of the RL agent, i.e., actor and critic networks, evaluation is then performed on all variants of both the training and test sets $\mathcal{T}^{\text{train}}\cup\mathcal{T}^{\text{test}}$. Figure˜3 illustrates how the environments are sampled in meta-training and adaptation phases. It should be noted that the RL updates need a separate buffer for keeping all the action and belief information in the interactions. To quantify adaptation to each task, the average return and success rate of all its variations are computed. This separation enables detailed analysis of generalization to familiar and unfamiliar environments under both identity and parametric diversity.
@@ 5.2 Training Implementation Details
    $\mathbf{u}_{t}^{s}=\text{FFN(Multi-Head(CausalSelfAttn}(f^{s}_{q}(\mathbf{E^{\prime}}^{s}),f^{s}_{k}(\mathbf{E^{\prime}}^{s}),f^{s}_{v}(\mathbf{E^{\prime}}^{s}))))$ ;
@@
    $\mathbf{u}_{t}^{r}=\text{FFN(Multi-Head(CausalSelfAttn}(f^{r}_{q}(\mathbf{E^{\prime}}^{r}),f^{r}_{k}(\mathbf{E^{\prime}}^{r}),f^{r}_{v}(\mathbf{E^{\prime}}^{r})))))$ ;
@@
    $h_{t}=\text{FFN(Multi-Head(CausalCrossAttn}(f^{r}_{q}(\mathbf{u}^{s}),f^{r}_{k}(\mathbf{u}^{s}),f^{r}_{v}(\mathbf{u}^{r})))))$ ;
@@ 5.3.1 Quantitative Findings
The quantitative evaluation focuses on two performance metrics, return and success rate, across all task variations in the MetaWorld ML-10 benchmark. The average values of these metrics are plotted against the number of training interaction frames in Figure˜4 and Figure˜5. Results are reported separately for training and test environments, ensuring that generalization can be meaningfully assessed. Table˜2 and Table˜3 present more detailed numerical findings, showing the post-training generalization performance in each environment for CRAFT and other Bayes-adaptive baselines.
@@
During many stages of training, CRAFT achieves higher final-episode return and success compared to the mean over the $n_{H}$ episodes in a meta-episode, denoted as $R_{f}-\overline{R}_{H^{+}}$ and $S_{f}-\overline{S}_{H^{+}}$, respectively. The corresponding values are plotted against the number of meta-training interaction frames in Figures˜6 and 7. As can be seen, this relative advantage remains in the later stages of training for the action-free methods, suggesting that the policy can still benefit from continued exploration and training. This indicates that accumulated experience within the meta-episode contributes meaningfully to improved performance, reflecting effective belief updates in the BAMDP framework. In contrast, other baseline methods show weaker improvement—or even degradation in some cases—between early and final episodes. The ability of transformer-based models to better extract information from long contexts compared to memory-based sequence models contributes to this result. It should also be noted that CRAFT, particularly when a state decoder is included during training, shows the most noticeable benefit during intermediate training stages, when the belief model is still developing and the policy is actively exploring. This further supports the value of gradient isolation in transformer-based context encoder training, which enables informative task inference independently and provides a strong prior for the adaptation process.
@@ 5.3.2 Qualitative Findings
To further analyze the structure of the learned latent space, the pairwise projections of the five-dimensional task belief vectors are plotted. A 2-D plot is generated for each pair of the latent space dimensions, resulting in ten distinct subplots covering all one-vs-one combinations shown in Figure˜8. Some task environments cluster closely in certain latent dimensions while remaining dispersed in others, suggesting that different latent axes are responsible for encoding distinct aspects of task variation. A 3-D projection of the latent vectors is illustrated in Figure˜9. By finding the means and variances of the 3-D UMAP projections computed for the latent representations of each of the 15 tasks [mcinnes2018umap], the estimated distributions are visualized as ellipsoids, indicating the variability region of each task in the latent space.
@@
Figure˜10 shows the average latent vector in all evaluation episodes of all variations of each environment. Some tasks consistently exhibit large values along specific latent dimensions, effectively dominating those axes. For instance, the lever-pull task is represented by a very low value along axis one and a large value on axis three, while the door-open task shows the highest value on axis one compared to all others. In general, conceptually similar tasks tend to share these dominant dimensions, indicating that the model has learned a structured and interpretable belief space. This organization reflects the transformer encoder-decoder’s ability to disentangle task-specific features. CRAFT was able to extract meaningful representations without access to action information, supporting its role as a reliable inference module in adaptive meta-RL.
Figure 10: Heatmap of mean task belief vectors inferred by CRAFT over all evaluation episodes in MetaWorld ML-10. Task-specific activation patterns and shared dominant latent dimensions among related tasks demonstrate disentangled representations aligned with semantics without action information.
@@ 7 Future Directions
The former possibility is quite straightforward in terms of implementation. However, the latter needs further explanation. As an example, consider a robotic environment expressed by two different MDPs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ per Equation˜4. These two MDPs share the same action space ($\mathcal{A}$), dynamic functions ($\mathcal{P}_{S},\mathcal{P}_{R}$), discount factor ($\gamma$), and episode length ($H$). However, the two MDPs differ in the state space domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ and optimal policies $\pi^{*}_{1}$ and $\pi^{*}_{2}$, respectively. For instance, $\mathcal{M}_{1}$ can receive state observations in the numerical space from sensor data, versus $\mathcal{M}_{2}$ receiving workspace images as observations. Assuming the essential Markov property holds for both MDPs, it is possible to infer a shared latent domain where $\Psi(\mathcal{S}_{1})\cong\Phi(\mathcal{S}_{2})$, as conceptualized in Figure˜11, such that $\Psi(\mathcal{M}_{1})\simeq\Phi(\mathcal{M}_{2})$. These transformed MDPs can be governed with a shared optimal policy $\pi^{*}\simeq\pi^{*}(\Psi(\cdot))\simeq\pi^{*}(\Phi(\cdot))$.
If learning the mappings $\Psi(\cdot)$ and $\Phi(\cdot)$, and their inverse functions $\Psi^{-1}(\cdot)$ and $\Phi^{-1}(\cdot)$, is tractable, it is possible to train an action-agnostic, and hence agent-agnostic, foundational transformer encoder-decoder from offline data in the source domain and fine-tune it with additional adapters to enable adaptation in the target domain. This means the proposed belief model can be trained on an MDP with numerical state observations. By learning observation adapters $q^{\text{belief}}$ and $q^{\text{policy}}$, and fine-tuning the RL agent consisting of value function $V_{\nu}$ and $\pi_{\zeta}$, the context-adaptive meta-RL pipeline can acclimate to visual state observations online as outlined in Figure˜12.